4x+5y=69 12x+7y=135

Answer by Boreal(15235)

4x+5y=69
12x+7y=135
multiply top by 3.
12x+15y=207
12x+7y=135
Now multiply top by (-1)
-12x-15y=-207
12x+7y =135 Add
-8y=-72
y=9. Put in first equation;;; 4x+ 45=69; 4x=24; x=6
Check second equation
12(6) +7(9)=72+63=135
(6,9)